OpenStudy (anonymous):

solve the equation using a u-substitution (x^2+x) ^2-8(x^2+x)+12=0

OpenStudy (anonymous):

Plug problem into mathway.com

OpenStudy (anonymous):

Let x^2 +x be a a^2 - 8a + 12 = 0

OpenStudy (anonymous):

factor that trinomial then substitute x^2+x back for a

OpenStudy (ash2326):

substitute u as x^+x we'll get u^2-8u+12=0 u^2-6u-2u+12=0 u(u-6)-2(u-6)=0 (u-2)(u-6)=0 u=2 or u=6 so x^2+x=2 and x^2+x=6 x^+x-2=0 (x+2)(x-1)-=0 x=-2 and x=1 from the other equation x^2+x-6=0 (x+3)(x-2)=0 x=-3 and x=2 so the solution for x=-3.-2.1 and 2

OpenStudy (kinggeorge):

First, let\[u=x^2+x\]so then we get\[u^2-8u+12=0 \longrightarrow (u-6)(u-2)=0\]so we have two possible answers\[\begin{matrix} u =6 \\ u=2 \end{matrix}\]Then we need to solve for x. Thus, if\[x^2+x=6\]then\[x^2+x-6=0 \longrightarrow (x+3)(x-2)=0\]so \[\begin{matrix} x=2 \\ x=-3 \end{matrix}\]Otherwise, \[x^2+x=2\]so\[x^2+x-2=0 \longrightarrow (x+2)(x-1)=0\]so\[\begin{matrix} x= -2 \\ x=1 \end{matrix}\]Overall, then, x has four possible values.\[\begin{matrix} x={-3} \\ x={-2} \\ x=1 \\ x=2 \end{matrix}\]

OpenStudy (anonymous):

thanks once again!!!

OpenStudy (kinggeorge):

Always glad to help.